Can someone help me on this problem?
I need to solve the following equation for x, and show proof.
3x - 5(3x -7) = 2(x+9) + 45
Thanks,
Sher

Found 2 solutions by jenrobrody, Earlsdon:
Answer by jenrobrody(19)

3x - 5(3x -7) = 2(x+9) + 45
Distribute: 3x -5(3x) -5(-7) = 2(x) + 2(9) +45
simplify: 3x - 15x + 35 = 2x +18 + 45
combine like terms: -12x + 35 = 2x + 63
add 12x to both sides: 35 = 2x + 12x + 63
combine like terms: 35 = 14x +63
subtract 63; 35 - 63 = 14x + 63 -63
simplify: -28 = 14x
divide by 14: -2 = x
check:
3(-2) - 5(3(-2)-7) = 2((-2)+9) + 45 ?
-6 - 5(-6 - 7) = 2( 7 ) + 45 ?
-6 - 5( -13) = 14 + 45 ?
-6 + 65 = 59 ?
59 = 59

Answer by Earlsdon(6103)

Solve for x:
3x-5(3x-7) = 2(x+9)+45 First, apply the distributive property on both sides.
3x-15x+35 = 2x+18+45 Now simplify this,
-12x+35 = 2x+63 Add 12x to both sides of the equation.
35 = 14x+63 Subtract 63 from both sides.
-28 = 14x Finally, divide both sides by 14.
-2 = x or x = -2